# Learn To be Efficient: Build Structured Sparsity in Large Language Models

### v1

@@ 4.2.2 Two-stage LTE Training
In practice, training the threshold-based Sigmoid router leads to three challenges. Firstly, Sigmoid function considers individual experts independently, making it hard to tease apart important experts from the rest. Secondly, the threshold-based indicator function in Equation 2 is non-differentiable, thereby preventing the training on routers. Lastly, manually choosing a threshold for each MoE layer can be impractical.
To tackle these challenges, we next propose a novel two-stage training algorithm with an efficiency loss penalty to MoEfy LLMs. As shown in Algorithm 1, our training algorithm consists of two training stages: a) model-router training stage and b) model adaption stage.
Stage 1: Model-router training. The goal of the model-router joint training stage is to train Sigmoid routers to capture the importance of experts for given inputs. In this stage, we jointly train routers and model parameters.
To address the non-differentiability issue, we switch the expert selection to a “soft” mode in this stage:
Instead of selecting a subset of experts, we always select all experts and multiply expert outputs with the corresponding expert score to make both router and model differentiable. Since there is no discrete selection in MoE layers, all parameters are trained for each input token.
@@
The efficiency loss calculates the mean of expert scores across all layers. This efficiency loss penalizes the output magnitudes of routers, driving routers to selectively allocate lower expert scores to less important experts, which helps distinguish experts with different importance. Moreover, instead of assigning the same sparsity for all layers, the efficiency loss also induces inter-layer competition, allowing LTE models to allocate adaptive sparsity for different layers. Our analysis in Figure 6 shows that LTE models are capable of adaptively allocating sparsity across different layers.
@@
where is the number of layers, and is a predefined threshold for all routers (we set for all experiments). The separability loss encourages router outputs to diverge from the threshold , which makes the router outputs more separable. An example is illustrated in Figure 3. The only difference between the models in the two figures is if the model is trained with the separability loss. We find that, when training without the separability loss, the expert scores are hard to separate, making it hard to decide the threshold. However, training with the separability loss markedly improves the separability of expert scores, allowing us to choose experts with the predefined threshold.
Combined with the task performance loss , our training loss in model-router training state is:
The efficiency coefficient, , is used to control the trade-off between inference efficiency and task performance. A higher indicates a larger inference cost penalty, leading to a more sparse model.
Different from the pretrained LLMs, MoE routers are randomly initiated. We use a larger learning rate (0.01) for routers’ parameters to make routers converge faster.
@@ 5 Experiment
In this section, we evaluate the effectiveness of LTE with different models on various datasets. In Section 5.2, the experiment results on NLU tasks show that LTE models achieve 80 - 95% sparsity in FFN layers without performance drop. Similarly, LTE also outperforms all other baselines on NLG tasks. For instance, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks.
@@ 5.1 Experimental Setting
Models. We implement LTE on both encoder-based models (RoBERTa, RoBERTa) and decoder-based models (GPT2-Medium, and LLaMA-7B) with the HuggingFace’s transformers. Following previous work (Zhang et al., 2022b), we set each expert to contain 32 neurons in FFN layers for all models.
Datasets. We evaluate LTE performance using 11 datasets across two setups:
Natural Language Understanding (NLU): we evaluate on eight tasks from the GLUE dataset (Wang et al., 2018) including sentiment classification:SST-2 (Socher et al., 2013), entailment: RTE (Dagan et al., 2005), grammatical classification: CoLA (Warstadt et al., 2018) two natural language reference tasks: MNLI (Williams et al., 2018), QNLI (Rajpurkar et al., 2016), and three tasks measuring sentence similarity: QQP (Iyer et al., 2017), STS-B (Cer et al., 2017), MPRC (Dolan & Brockett, 2005). We report the Pearson correlation coefficient for STS-B, the Matthews correlation coefficient for CoLA, and the accuracy for the other six datasets.
Natural Language Generation (NLG): we evaluate on data-to-text (E2E Challenge (Novikova et al., 2017)), text-to-text summarization (XSum (Narayan et al., 2018)), and next token prediction language modeling (Wikitext103 (Merity et al., 2016)). For both E2E and XSum, we report the ROUGE-L (Lin, 2004) to measure the generation text quality and report the perplexity performance for the Wikitext dataset.
Baselines. (1) MoEfication (Zhang et al., 2022b; Liu et al., 2023): MoEfication proposes to replace soft activations with ReLU and then fine-tune the model on downstream datasets. In our evaluation, we report the number of MoEfication with parameter clustering and ground-truth expert selection. Note that ground-truth selection is not practically feasible but can be considered as an upper bound for expert selection. (2) Keep-the-Largest-Activation (KLA): To explore the sparsity upper-bound in LLMs trained with task performance loss alone, we manually create an upper-bound baseline based on an ideal scenario. We assume that an oracle can always predict the activation value of each neuron with no cost. Given a sparsity level, the oracle always selects neurons with the largest absolute activation values. Although KLA is not feasible in real-world settings, it represents an upper-bound performance for models that do not optimize efficiency in training. The models used in KLA are also models fine-tuned on downstream tasks.
5.2 Performance Comparison
@@
No performance drop with 80-95% sparsity on NLU tasks. We first evaluate our methods on eight NLU Tasks in GLUE datasets(Wang et al., 2018) with two encoder-based models (RoBERTa and RoBERTa). Due to the space limit, we report four datasets (MRPC, SST2, QNLI, and MNLI) in Figure 4 and other four datasets (RTE, CoLA, STSB, and QQP) in Figure 8 in Appendix B.
As shown in Figure 4, our method (LTE) consistently outperforms the baseline models on all datasets. Both MoEfication and KLA exhibit performance drop when sparsity reaches 70%–80%, while LTE maintains good performance even at higher sparsity levels (90%). The major reason behind this performance gap is that, although LLMs naturally have high sparsity, they are still not optimized for efficiency. The efficiency loss penalty in LTE training loss effectively enables LLMs to activate fewer neurons while minimally impacting task performance.
@@
Moreover, we notice that while ReLU-based models achieve high sparsity with a slight decrease in performance, replacing GeLU with ReLU in RoBERTa can affect performance on some datasets. For instance, replacing GeLU with ReLU in a RoBERTa model results in a 10% accuracy drop on the MRPC dataset.
@@
50% FLOPs saving on NLG tasks. We next evaluate LTE with decoder-based models on three types of generation tasks. Similar to NLU tasks, Figure 5 shows that our method outperforms all other baselines. NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level compared to NLU tasks. However, LTE-trained models have better resilience to the sparsity increase. Especially at a high sparsity level, other baselines fail to generate meaningful content, but LTE still generates content with a marginal quality drop. In NLG tasks, we also observe that replacing soft activation with ReLU can hurt model performance. However, training with ReLU-LLaMA largely mitigates some of the performance drops.
| XSum | E2E | WikiText-103 | |
@@
To better understand the computation saving of LTE, we compare the FLOPs per token of different methods on LLaMA, allowing an up to 0.05 ROUGE-L decrease for XSum and E2E, and an up to 0.5 perplexity (PPL) increase for WikiText-103. It is noteworthy that the FLOPs for routers are included in our method, constituting approximately 1% of the total FLOPs of FFN layers. The results are presented in Table 1. The evaluation results show that LTE provides a 1.83x - 2.59x FLOPs speed-up and gives the most FLOPs savings among all methods. Those evaluation results demonstrate that LTE effectively provides higher sparsity for LLMs even with soft activation functions.
5.3 Ablation Study & Analysis
Another intriguing finding is that RoBERTa and GPT2-M have very different sparsity patterns. In RoBERTa, the sparsity of MoE layers significantly increases with depth, contrasting with GPT2-M, where MoE layer sparsity decreases as layer depth increases. This difference can stem from differences in the design of these two transformers and tasks. Given that MRPC is a sentence classification task, encoder-based transformers like RoBERTa primarily rely on the ”[CLS]” token for classification. Consequently, other tokens become less important in deeper layers, allowing sparser FFN layers. In contrast, GPT2, a decoder-based transformer designed for language generation, requires all tokens to generate the next token. Furthermore, recent research (Geva et al., 2021) on interpretability suggests that deeper FFN layers store more complex semantic information, which may drive deep FFN layers to be less sparse.
@@ 6 Conclusion
Appendix A Experiment Settings
We presented all training hyperparameters in Table 2, 3, 4, and 5. For hyperparameters presented in {}, we select the best hyperparameter for each task. We follow the settings in RoBERTa paper (Liu et al., 2019) to fine-tune RoBERTa on GLUE datasets. We set the coefficient for the separability loss ( in Equation 6) to be 0.5 for all stage 1 training.
Appendix B Additional Evaluation Results

### v3

@@ 4.2.2 Two-stage LTE Training
Training the router poses three practical challenges. Firstly, as we consider training experts independently to mitigate bias (e.g., using a Sigmoid routing function), it is difficult to tease apart important experts from the rest. Secondly, the non-differentiability of threshold-based expert selection further complicates router training. Lastly, setting thresholds for each MoE layer is non-trivial.
To tackle these challenges, we next propose a novel two-stage training algorithm with an efficiency loss penalty to MoEfy LLMs. Our training algorithm consists of two training stages: a) model-router training stage and b) model adaption stage.
Stage 1: Model-router training. In this stage, we jointly train routers and model parameters to capture the importance of experts for given inputs in the Sigmoid routers. To address the non-differentiability issue, we switch the expert selection to a “soft” mode:
Instead of selecting a subset of experts, we always select all experts and multiply expert outputs with the corresponding expert score to make both router and model differentiable. Since there is no discrete selection in MoE layers, all parameters are trained for each iteration.
@@
where is the number of layers, and is the number of experts in each layer. The efficiency loss calculates the mean of expert scores across all layers. This efficiency loss penalizes the output magnitudes of routers, driving routers to selectively allocate lower expert scores to less important experts, which helps distinguish experts with different importance. Moreover, instead of assigning the same sparsity for all layers, the efficiency loss also induces inter-layer orchestration, allowing LTE-trained models to allocate adaptive sparsity for different layers. We show that LTE models are capable of adaptively allocating sparsity across different layers (Figure 15).
@@
where is a predefined threshold for all routers (we set for all experiments). The separability loss encourages router outputs to diverge from the threshold , which makes the router outputs more separable. An example is illustrated in Figure 3. The only difference between the models in the two figures is if the model is trained with a separability loss. We find that the expert scores are not separable when training without the separability loss, making it hard to decide the threshold. However, training with the separability loss markedly improves the separability of expert scores, allowing us to choose experts with the predefined threshold.
Combined with the task performance loss , our training loss in model-router training stage is:
The efficiency coefficient, , is used to control the trade-off between inference efficiency and task performance, and can be used to control the sparsity of trained models. A higher indicates a larger inference cost penalty, leading to a more sparse model. is the separability loss, we set for our evaluation. We conduct an ablation study on both hyperparameters in Section 5.4.
@@
4.3 Hardware-efficient Implementation
As illustrated in Figure 4, same as Deja Vu [20] and moefication [43], LTE provides very structured sparsity. After selecting neurons in intermediate layers, a CUDA kernel needs only to load the relevant columns and rows of and from GPU global memory to SRAM to compute dot product, thus reducing memory-I/O as well as computational overheads. Thanks to the structured sparsity provided by LTE, the kernel avoids non-coalesced memory access by storing column-major and row-major . In this paper, we use Triton 2.3.0 to implement a customized MLP layer to translate the sparsity to wall-clock time latency reduction. The evaluation of our custom kernel is presented in Section 5.3.
@@ 5 Experiment
In this section, we conduct extensive evaluationS to verify the effectiveness of LTE. For instance, in Section 5.2, LLaMA with LTE provides a 1.83x - 2.59x FLOPs speed-up on NLG tasks. Along with our hardware-aware custom kernel, LTE reduces 25% LLaMA2-7B inference latency at 50% sparsity.
@@ 5.1 Experimental Setting
Models. We implement LTE on both encoder-based models (RoBERTa, RoBERTa) and decoder-based models (GPT2-Medium, and LLaMA-7B) with the HuggingFace’s transformers. Following previous work [43], we set each expert to contain 32 neurons in FFN layers for all models.
Datasets: 1) Natural Language Understanding (NLU): we evaluate on eight tasks from the GLUE dataset [37] SST-2 [32], RTE [4], CoLA [39], MNLI [40], QNLI [28], QQP [12], STS-B [3], and MPRC [5]. We report the Pearson correlation coefficient for STS-B, the Matthews correlation coefficient for CoLA, and the accuracy for the other six datasets. 2) Natural Language Generation (NLG): we evaluate LTE on E2E [26], XSum [25], and Wikitext103 [23]. For both E2E and XSum, we report the ROUGE-L [18] to measure the generation text quality and report the perplexity performance for the Wikitext dataset. 3) Instruction Tuning: besides downstream tasks, we also evaluate LTE on instruction tuning tasks to evaluate LTE’s generalization capabilities. We use Tulu dataset [38] to perform instruction tunning with LTE, and we evaluate models with MMLU benchmark [10]). MMLU benchmark is a comprehensive evaluation dataset designed to evaluate the generalization capabilities of LLMs and consists of 15908 questions from 57 distinct tasks.
Baselines. We compare two baselines with LTE in our paper: (1) MoEfication [43]: MoEfication proposes to replace soft activations with ReLU and then fine-tune the model on downstream datasets. In our evaluation, we report the number of MoEfication with parameter clustering and ground-truth expert selection. (2) Deja Vu [20]: Deja Vu employs predictors to estimate the values of neurons in intermediate layers, pruning neurons with absolute magnitudes (absolute values for non-ReLU activations). The models used in Deja Vu evaluation are models fine-tuned on specific tasks as well.
5.2 End-to-end Performance Comparison
LTE, MoEfication, and Deja Vu provide the same type of structured sparsity in FFN layers. For the convenience of comparison, we use FFN neuron sparsity as a measure of efficiency in this section. We will further discuss how our hardware-efficient implementation translates this sparsity into wall-clock time latency reduction in Section 5.3. For our method (LTE), we get models with different sparsity by tuning efficiency loss in Equation 6.
@@ 5.2 End-to-end Performance Comparison
No performance drop with 80-95% sparsity on NLU tasks. We first evaluate our methods on eight NLU Tasks in GLUE datasets[37] with two encoder-based models (RoBERTa and RoBERTa). Due to the space limit, we report four datasets (MRPC, SST2, QNLI, and MNLI) on RoBERTa in Figure 5 and other results in Figure 13 in Appendix C.
@@
As shown in Figure 5 and Figure 13, our method (LTE) consistently outperforms the baselines on all datasets. MoEfication exhibits a performance drop when sparsity reaches 70%–80%, and Deja Vu experiences performance drop at a relatively low sparsity level (e.g., 30%), while LTE maintains good performance even at higher sparsity levels (90%). Moreover, we notice that while ReLU-based models achieve high sparsity with a slight decrease in performance, replacing GeLU with ReLU in RoBERTa can affect performance on some datasets. For instance, replacing GeLU with ReLU in a RoBERTa model results in around 10% accuracy drop on the MRPC dataset.
@@
50% FLOPs saving on NLG tasks. We next evaluate LTE with decoder-based models on three types of generation tasks. Similar to NLU tasks, Figure 6 shows that LTE outperforms all other baselines. NLG tasks are more challenging: all methods start to have a performance drop at a lower sparsity level
| XSum | E2E | Wiki | |
@@
compared to NLU tasks. Yet, LTE-trained models have better resilience to the sparsity increase. Especially at a high sparsity level, while other baselines fail to generate meaningful content, LTE still generates content with a marginal quality drop. Similarly, we also observe the performance drop caused by replacing soft activation with ReLU.
@@
Additionally, to better understand the computation saving of LTE, we compare the FLOPs per token of different methods on LLaMA-7B, allowing an up to 0.05 ROUGE-L decrease for XSum and E2E, and an up to 0.5 perplexity (PPL) increase for WikiText-103. It is noteworthy that the FLOPs for routers are included in our method, constituting approximately 1% of the total FLOPs of FFN layers. The results are presented in Table 1. The evaluation results show that LTE provides a 1.83x - 2.59x FLOPs speed-up and gives the most FLOPs savings among all methods.
LTE Outperforms SOTA on instruction tuning. To further evaluate the effectiveness of LTE, we evaluate LTE as a chatbot and verify model performance in a few-shot learning scenario. We fine-tune LTE on the Tulu instruction tunning dataset [38], and then evaluate the few-shot performance on MMLU benchmark [10]. We use LLaMA2-7B as the base model for LTE training. For a fair comparison, we also fine-tune base models with the Tulu dataset for other baselines. Specifically, for Deja Vu, we use Tulu-fine-tuned LLaMA2-7B as the base model. For Moefication, we use Tulu-fine-tuned ReLULLaMA-7B as the base model to ensure the best possible baseline performance. We present the 5-shot MMLU accuracy comparison in Figure 7. The evaluation results show that LTE outperforms other baselines across all sparsity levels. Yet, LTE shows a weaker sparsity-performance trade-off compared to downstream tasks. Our hypothesis is that LTE routers change the structure of the original LLaMA model, which influences general language patterns learned by models in pretraining. Instruction tuning alone is not enough to fully recover this ability (considering instruction tuning has much smaller datasets). We believe that training with a wider variety of data (like pretraining data) will further improve LTE performance.
5.3 Translate Sparsity to Wall-clock Time Speedup
25% wall-clock time latency reduction. As discussed in Section 4.3, the sparsity introduced by LTE is highly structured and can be easily translated to latency reduction. This section demonstrates that our custom Triton kernel effectively translates the LTE sparsity to wall-clock time speed up. We evaluate wall-clock time speed up using LLaMA2-7B on a single 3090Ti. We also reported vanilla Pytorch indexed matrix multiplication. Due to the requirement for new memory allocations, slicing in Pytorch is very time-consuming. This makes indexed matrix multiplication even slower than dense matrix multiplication at most sparsity levels. We report the wall-clock time latency comparison in Figure 8 (noting that the router inference latency is already included in LTE latency). Compared to a dense FFN block, our custom Triton kernel achieves nearly linear speed-up with respect to sparsity (Left). At around 50% sparsity, LTE reduces end-to-end generation latency by approximately 25%.
5.4 Ablation Study
Experts grouping algorithms. Three different expert grouping algorithms are explored in [43]: random grouping, parameter K-means clustering, and co-activation graph split. The co-activation graph split algorithm constructs a graph where each neuron represents a node and the edges denote the frequency of co-activation between neuron pairs. This graph is subsequently divided into subgraphs using the graph split algorithm [15], with each subgraph representing a group of neurons. To study the effectiveness of these grouping algorithms, we compare their LTE performance by fine-tuning GPT2 on the WikiText-103 dataset. As shown in Figure 11, both co-activation graph split and parameter K-means grouping have very similar performance (similar to finding in [43]) and outperform random grouping. Given that parameter K-means grouping is simpler to implement, as co-activation graph split requires collecting activation data, we use K-means grouping in our paper.
The relation between efficiency loss hyperparameter and sparsity. As discussed in Section 4.2.2, efficiency loss hyperparameter is used to balance the trade-off between task performance and sparsity (efficiency). We illustrate the relation between and sparsity across different models and tasks in Figure 11. For a given task and model, increasing results in greater sparsity.
Ablation study on separability loss. Another hyperparameter in LTE training, , is used to encourage the separability of router outputs, which facilitates us to choose the threshold to pick neurons (as shown in Figure 3). We compare models trained with different lambda in Figure 11. When increasing from (i.e., no separability loss) to , we observe a constant perplexity drop for all sparsity. However, if we keep increasing , perplexity does not further decrease.
@@ 6 Conclusion
Limitations
While LTE outperforms other baselines, LTE needs further fine-tuning to get contextual sparse models. (Yet, training is a one-time investment that yields long-term benefits for LLM serving.) A future study is to combine LTE with other parameter-efficient fine-tuning techniques to reduce LTE training overhead. Another limitation is our computational resource, which prevents us from training LTE on pretraining or RLHF. As discussed in Section 5.2, we believe that training with a wider variety of data will further improve LTE performance. We leave this for future exploration.
Appendix A Code and Datasets License
Codebase. Our model implementation is based on Huggingface transformer repo: Apache-2.0 license
Datasets. We list the license of used datasets as follow:
GLUE dataset [37]: Custom License;
WikiText103 [23]: CC BY-SA 3.0;
XSum [25]: MIT License;
E2E [26]: CC4.0-BY-SA;
MMLU [10]: Custom License.
Appendix B Experiment Settings
We presented all training hyperparameters in Table 2, 3, 4, and 5. For hyperparameters presented in {}, we select the best hyperparameter for each task. We follow the settings in RoBERTa paper [19] to fine-tune RoBERTa on GLUE datasets. We set the coefficient for the separability loss ( in Equation 6) to be 0.5 for all stage 1 training. We use different to control the sparsity of trained models (Figure 11). Hardware: All models are trained and evaluated on A100, A40, and 3090Ti, depending on memory usage and availability.
Appendix C Additional Evaluation Results
C.1 NLU Performance Comparison
@@
C.2 Further Analysis on Sparsity Provided by LTE
Contextual Sparsity for Larger Batches. Similar to previous contextual sparsity work [20], we observe that as the batch size increases, more neurons are activated, leading to a reduction in union sparsity. These dynamics are illustrated in Figure14. Similar to Deja Vu [20], the number of activated neurons does not grow linearly with the batch size, indicating a non-uniform distribution of parameter access from different inputs. This property can extend LTE for a high-throughput batch setting by batching inputs activating similar neurons together to achieve a high union sparsity.
Another intriguing finding is that RoBERTa and GPT2-M have very different sparsity patterns. In RoBERTa, the sparsity of MoE layers significantly increases with depth, contrasting with GPT2-M, where MoE layer sparsity decreases as layer depth increases. This difference can stem from differences in the design of these two transformers and tasks. Given that MRPC is a sentence classification task, encoder-based transformers like RoBERTa primarily rely on the "[CLS]" token for classification. Consequently, other tokens become less important in deeper layers, allowing sparser FFN layers. In contrast, GPT2, a decoder-based transformer designed for language generation, requires all tokens to generate the next token. Furthermore, recent research [9] on interpretability suggests that deeper FFN layers store more complex semantic information, which may drive deep FFN layers to be less sparse.
@@ C.2 Further Analysis on Sparsity Provided by LTE
| Hyperparameter | Fine-tuning | LTE-Stage 1 | LTE-Stage 2 |
| Learning rate | 2e-5 | 2e-5 | 2e-5 |
| Training batch size | 128 | 128 | 128 |
| Training epochs | 3 | 1 | 4 |
| Weight decay | 0 | 0 | 0 |
| Warm up ratio | 0.06 | 0.06 | 0.06 |